OpenStudy (anonymous):

Please help me p(x)=x^3+7x^2+4x x =-2

OpenStudy (anonymous):

I think it is asking to evaluate the function at x=-2, so you plug it in \(p(-2)=(-2)^3+7(-2)^2+4(-2)\) and solve \(p(-2)=-8+7(4)-8\) \(p(-2)=28-16\) \(p(-2)=12\)

OpenStudy (anonymous):

thank you so much but it says using the remainder therorm so there si another part

OpenStudy (anonymous):

Okay this is different then, instead of direct substitution you will use division, have you ever heard of synthetic division?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay here is how you do this problem: